# Performance of Practical Quantum Oblivious Key Distribution

### v2

@@ I Introduction
In the context of quantum cryptography, OT is remarkable because, unlike classically, there exists a reduction from OT to commitment schemes [21, 22, 23]. This result is somewhat undermined by the existence of several theorems regarding the impossibility of unconditionally secure commitments both in classical [24] and quantum [25, 26] cryptography, and it was further proven impossible in the more general abstract cryptography framework [27]. These results, in turn, imply that unconditionally secure OT itself is impossible. In light of this, approaches with different technological or physical constraints on the adversarial power have been proposed. Practical solutions based on hardware limitations, such as bounded and noisy storage [28, 29, 30, 31], have the disadvantage that the performance of such protocols decreases as technology improves.
@@
The contributions of this work can be summarized as follows: We introduce a definition for a quantum ROT protocol, satisfying a strong indistinguishability-based security notion, which generalizes the security of classical ROT protocols. We present a protocol that realizes said quantum ROT, based on the standard commitment-based construction (originally by [41, 21]). The protocol uses a weakly-interactive string commitment scheme which is statistically binding and computationally hiding, and can be implemented in practice using current QKD setups. We present a formal finite-key security proof of the proposed protocol assuming only the existence of quantum-secure OWFs, together with security bounds as functions of the protocol’s parameters. We also present calculations for the maximum usable channel error, as well as for the key rate as a function of the number of shared signals per instance of the protocol. We experimentally demonstrate our protocol using current technology with a setup based on polarization-entangled photons. We also present a security analysis which accounts for potential implementation-specific attacks and how they can be circumvented using an appropriate reporting strategy. Finally, we compare our performance results with the performance of current ROT solutions and point out the advantages and disadvantages of using quantum ROT in the context of MPC.
II Preliminaries
II.1 Quantum computational efficiency and distinguishability
A quantum program is defined as a set of instructions to be run in a specified type of quantum computer, which are devices that have a classical interface and a quantum part. Quantum programs can be compared with probabilistic classical programs as they both have natural numbers as inputs/outputs. When a quantum computer runs the program $T$ with input $x\in\mathbb{N}$, we assume that the quantum part of the computer starts with some predefined initial state, performs a sequence of operations on its quantum registers, and upon halting, it outputs $T(x)\in\mathbb{N}$ on its classical interface by reading the appropriate registers associated with the program’s output.
Definition II.1.
(Computational efficiency)
Definition II.2.
(Distinguishing Advantage)
Definition II.3.
(Indistinguishability - Finite)
Proposition II.1.
$\rho_{1}\approx_{D(\rho_{1},\rho_{2})}\rho_{2}.$ (4)
Definition II.4.
(Indistinguishability – Asymptotical)
For the purposes of this work, the concept of indistinguishability will be often used to compare the state of systems in a “real” run of the protocol versus another “ideal” desired state. In those cases, the associated families of states are parameterized by the protocol’s security parameters. When the associated families $\{\rho_{1}^{(k)}\}$ and $\{\rho_{2}^{(k)}\}$ are clear by context, we will just refer to them as $\rho_{1}$ and $\rho_{2}$ and use the following notation to denote indistinguishability:
[…]
Definition II.5.
(Efficient quantum operation)
The following properties are straightforward to prove from Definitions II.3 and II.4 and the properties of trace distance:
Lemma II.1.
(Properties of indistinguishable states I)
Let $\rho_{1},\rho_{2},\rho_{3}\in\mathcal{D}(\mathcal{H})$:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
Lemma II.2.
(Properties of indistinguishable states II)
  1. 1.
  2. 2.
  3. 3.
  4. 4.
II.2 Entropic quantities
Definition II.6.
(Max-entropy)
Definition II.7.
(Min-entropy and conditional min-entropy)
$H_{\textnormal{min}}(\rho)=-\log(\lambda_{\max}(\rho)).$ (8)
Definition II.8.
(Partially classical states)
Lemma II.3.
(Properties of min- and max-entropy)
Let $\varepsilon,\varepsilon^{\prime}\geq 0$:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
We use universal hashing to implement randomness extraction in the final steps of the protocol. The proof both Lemmas II.3 and II.4 can be found in [44].
Definition II.9.
(Universal hashing)
$\textnormal{Pr}[f_{i}(x)=f_{i}(y)]\leq\frac{1}{2^{n}}.$ (13)
Lemma II.4.
(Quantum leftover hash)
$\otimes|x\rangle\ \!\!\!\langle x|_{A}\rho^{x}_{B}\Big{)},$ (14)
it holds that
with
II.3 Quantum Random Oblivious Transfer (ROT)
Definition II.10.
@@ III The protocol
In order to facilitate the finite-key security analysis, the description of $\pi_{\textnormal{QROT}}$ features two statistical tolerance parameters, denoted as $\delta_{1},\delta_{2}$. The role of $\delta_{1}$ is to account for the error in the estimation of the Qubit Error Rate (QBER), while the role of $\delta_{2}$ is to account for the small variations in the frequency of outcomes of $/50$ 50 events. These parameters can be ignored (set to zero) when considering very large values of $N_{0}$. In the following description of the protocol we use the bit values $0,1$ to the denote the computational and Hadamard bases for qubit Hilbert spaces, respectively, and ${|{\Phi^{+}}\rangle}=\frac{1}{\sqrt{2}}({|{00}\rangle}+{|{11}\rangle})$. Finally, we will use the relative (or normalized) Hamming weight function $r_{H}:\{0,1\}^{n}\rightarrow[0,1]$ defined for any $x=(x_{1},\ldots,x_{n})$ as
@@
  • Parameter estimation sample ratio $10<\alpha<$ 1
@@
  • Maximum qubit error rate $20\leq p_{\text{max}}\leq 1/$ 2
@@
III.1 Security and performance of the main protocol
@@
A high-level proof of Theorem III.1, including the derivation of the security bounds from Lemmas III.1 and III.2 can be found in Section IV and further details can be found in Appendix A. The security of $\pi_{\textnormal{QROT}}$ is given by its main security parameter $N_{0}$, as well as the security parameters of the underlying commitment and IR schemes $k$ and $k^{\prime}$, respectively. These values can be computed for the statistical security features of the protocol and are given by the following lemmas:
Lemma III.1.
@@
For the purposes of this analysis, we assume that the commitment and IR schemes, as well as their security parameters $k,k^{\prime}$, are appropriately chosen to satisfy the desired security level and we focus in the dependence $\varepsilon_{\max}$ has on the channel losses, characterized by the parameter $p_{\max}$, and the number of quantum signals $N_{0}$. We are also interested in a quantity known as the secret key rate $R_{\textnormal{key}}$. For given values of $N_{0}$, $\alpha$, $\delta_{1}$, $\delta_{2}$, $p_{\max}$, and $\varepsilon_{\max}$, let $n_{\max}$ be the largest number for which the associated $n_{\max}$-bit ROT has at least security $\varepsilon_{\max}$, then
@@
Figure 1: Maximum key rate output $\frac{n}{N_{0}}$ versus error rate $p_{\max}$. The blue line represents the upper bound for the key rate, when $N_{0}\rightarrow\infty$, $\alpha,\delta_{1},\delta_{2}$ are taken to be 0 and $f=1$. The orange line represent a more typical case with $\alpha=0.35$, $\delta_{1}=0.01,\delta_{2}=0.025$, and $f=1.2$.
Figure 2: Maximum key rate behaviour as a function of $N_{0}$ for different security levels.
@@
III.2 Experimental implementation performance
@@
where $N_{\textnormal{crit}}$ is computed using the optimal values of $\alpha,\delta_{1},\delta_{2}$ for the respective error rate $p_{\max}$ and undetected multi-photon rate $p_{\textnormal{multi}}$ associated to $P$, assuming perfectly efficient information reconciliation, $f=1$ (see Section V for the details on the implementation and its security). As seen in Figure 4, for this implementation, the additional coincidence rate gained by increasing $P$ is not enough to compensate for the induced increased error. This result is not immediately obvious, as $N_{\textnormal{crit}}$ does not depend explicitly on $p_{\max}$. The decrease in performance comes from the fact that increasing $p_{\max}$ limits the values that $\delta_{1}$ can have while maintaining positive key rates. This restriction on the values of $\delta_{1}$ ultimately results in an increase in $N_{\textnormal{crit}}$ and therefore, a reduction on $R_{OT}$.
@@
Table 1 shows the an example of the performance of the protocol in a real-world implementation using the data from the experimental setup. For the commit/open phase, the weakly-interactive string commitment protocol introduced in [37] was implemented using the BLAKE3 hash function algorithm as a one-way function. For the post-processing phase, a low density parity check (LDPC) code was used for IR, and random binary matrices were used to implement the universal hash family for privacy amplification. We evaluated the performance by the number of $128$-bit ROT instances able to be completed per second (It is worth noting that, using a Mac mini M1 2020 16GB computer, the post-processing throughput was enough to handle all the data from the experiment, the bottleneck being the quantum signal generation rate).
@@
[…]
@@
[…]
@@
[…]
@@
IV Security Analysis
In this section, we prove the main security result, which relates the overall security of the protocol as a function of its parameters $N_{0},\alpha,\delta_{1}$, and $\delta_{2}$ in Theorem III.1. For clarity of presentation, we have compacted some of the properties into lemmas, for which detailed proofs can be found in Appendix A.
IV.1 Correctness
@@
Recall that we model the aborted state of the protocol as the zero operator. This way, whenever we have a mixture of states, some of which trigger aborting and some that do not, the abort operation removes the events that trigger it from the mixture, effectively reducing its trace by the probability of aborting. There are three instances where the protocol can abort: first during Step (7) if the estimated qubit error rate is larger than $p_{\max}$; the second one is during Step (9) if Bob does not have enough (mis)matching bases to construct the sets $I_{0},I_{1}$; and finally during Step (11) if the IR verification fails. The probability of aborting in Steps (7) and (11) depends on the particular transformation that the states undergo when being sent from Alice’s to Bob’s laboratory, about which we make no assumptions. We can group these three abort events and denote by $p_{\textnormal{abort}}$ the probability of the protocol aborting by the end of Step (11). The state at this point can be written as $(1-p_{\textnormal{abort}})\rho^{\top}$, where $\rho^{\top}$ represents the normalized state conditioned that the protocol has not aborted by this point. As Lemma IV.1 states, the verifiability property of the Information Reconciliation scheme guarantees that the states that “survive” past Step (11) have the property that Bob’s corrected string $y^{B}$ is the same as Alice’s outcome string $x^{A}_{I_{0}}$, which is uniformly distributed.
Lemma IV.1.
@@
During Step (12) universal hashing is used in both $x^{A}_{I_{0}}$ and $x^{A}_{I_{1}}$ to obtain $m_{c}$ and $m_{\bar{c}}$. Because Eq. (34) describes a state for which the $X^{A}_{I_{0}},X^{A}_{I_{1}}$, and $C$ subsystems are independent and uniformly distributed, it follows from Lemma IV.1 that
@@
Finally, using the quantum leftover hash Lemma II.4 twice (once for $m_{0}$ and $m_{1}$) with the corresponding entropy terms given by Eq. (IV.1), together with Lemma II.1 (1), we conclude that the state $\rho^{(\textnormal{out})}_{M_{0},M_{1},C,M_{C}}$ of the output systems after the post processing phase satisfies (substituting $p_{\textnormal{succ}}=1-p_{\textnormal{abort}}$)
@@
IV.2 Security against dishonest sender

### v3

@@ 1 Introduction
In the context of quantum cryptography, OT is remarkable because, unlike classically, there exists a reduction from OT to commitment schemes [21]. This result is somewhat undermined by the existence of several theorems regarding the impossibility of unconditionally secure commitments both in classical [22] and quantum [23, 24] cryptography, and it was further proven impossible in the more general abstract cryptography framework [25]. These results, in turn, imply that unconditionally secure OT itself is impossible. In light of this, approaches with different technological or physical constraints on the adversarial power have been proposed. Practical solutions based on hardware limitations, such as bounded and noisy storage [26, 27, 28, 29], have the disadvantage that the performance of such protocols decreases as technology improves.
@@
The idea of using quantum conjugate coding and commitments for oblivious transfer was originally proposed by Crépeau and Kilian [17] and then refined by Bennet et al, in [21] with the BBCS92 protocol (shown in Fig. 1). This construction has been extensively studied from the point of view of its theoretical security [39, 40, 41, 38, 42, 43]. However, while practical security analyses and experimental implementations have been made for quantum OT in the noisy storage model [28, 29], there are no works analyzing the quantum resource requirements and the resulting performance of implementing the BBCS92 protocol using existing computationally-secure commitment schemes based on OWFs. Such analyses are needed to demonstrate secure experimental implementations, and provide an important step in bringing quantum OT to real-world usage.
BBCS92 Quantum OT protocol Parties: The sender Alice and the receiver Bob.
1. Alice prepares $N$ entangled states of the form $\frac{1}{\sqrt{2}}({|{00}\rangle}+{|{11}\rangle})$ and, for each state prepared, sends one of the qubits to Bob. 2. Alice randomly chooses a measurement bases string $\theta^{A}\in\{+,\times\}^{N}$ and, for each $i=1,\ldots,N$ measures her share of the $i$-th entangled state in the $\theta_{i}$ basis to obtain outcome $x^{A}_{i}$ and the outcome string $x^{A}=(x^{A}_{1},\ldots,x^{A}_{N})$. 3. Bob uses the same process to obtain the measurement bases and outcome strings $\theta^{B}$ and $x^{B}$, respectively. 4. For each $i$, Bob commits $(\theta^{B}_{i},x^{B}_{i})$ to Alice. 5. Alice chooses randomly a set of indices $T\subset\{1,\ldots,N\}$ of some fixed size and sends $T$ to Bob. 6. For each $j\in T$, Bob opens the commitments associated to $(\theta^{B}_{j},x^{B}_{j})$. 7. Alice checks that $x^{A}_{j}=x^{B}_{j}$ whenever $\theta^{A}_{j}=\theta^{B}_{j}$ within the test set. If the test fails Alice aborts the protocol, otherwise she sends the string $\theta^{A}$ to Bob. 8. Bob separates the remaining indices in two sets: $I_{0}$ - the indices where Bob’s measurement bases match Alice’s, and $I_{1}$ - the set of indices where their bases do not match. Then, he samples randomly $c$ and sends the ordered pair $(I_{c},I_{\bar{c}})$ to Alice. 9. Alice defines the strings $\mathbf{x}^{A}_{c},\mathbf{x}^{A}_{\bar{c}}$ using the indices in the respective sets $(I_{c},I_{\bar{c}})$. Then, she samples randomly a function $f$ from a universal hash family, sends $f$ to Bob and outputs $m_{c}=f(\mathbf{x}_{0})$ and $m_{1}=f(\mathbf{x}_{\bar{c}})$ to Bob. 10. Similarly, Bob defines the string $\mathbf{x}^{B}$ from the set $I_{0}$ and outputs $m_{c}=f(\mathbf{x}^{B})$ and $c$.
Figure 1: Quantum oblivious transfer protocol based on commitments
Motivated by practical considerations, we consider Naor-style statistically binding and computationally hiding commitments, as these are well understood and efficient to implement (note that stronger commitments can be considered, such as the quantum-based commitments studied in [38, 42], however, implementing those requires significantly more computational and quantum resources).
The contributions of this work can be summarized as follows:
We introduce the definition for a quantum ROT protocol, satisfying a strong indistinguishability-based security notion equivalent to the one presented in [44], which generalizes the security of classical ROT protocols. We present a protocol that realizes said quantum ROT based on the BBCS construction. The protocol uses a weakly-interactive string commitment scheme which is statistically binding and computationally hiding, and can be implemented in practice using current QKD setups.
@@
We present a formal finite-key security proof of the proposed protocol accounting for noisy quantum channels assuming only the existence of quantum-secure OWFs, together with security bounds as functions of the protocol’s parameters. We also present calculations for the maximum usable channel error, as well as for the key rate as a function of the number of shared signals per instance of the protocol. Additionally, we study the composability properties of said protocol. In particular, we show that there is a family of weakly-interactive commitments which, when used in the quantum OT protocol, result in universally composable quantum OT in the classical access random oracle model. We experimentally demonstrate our protocol using current technology with a setup based on polarization-entangled photons. We also present a security analysis which accounts for potential implementation-specific attacks and how they can be circumvented using an appropriate reporting strategy. Finally, we compare our performance results with the performance of current ROT solutions and point out the advantages and disadvantages of using quantum ROT in the context of MPC.
2 Quantum Random Oblivious Transfer (ROT)
In this work, the concept of indistinguishability will be often used to compare the state of systems in a “real” run of the protocol versus another “ideal” desired state. These relations are defined over families of quantum states parametrized by the security parameter of the respective protocol. Hence, indistinguishability relations are statements on the asymptotic behavior of the protocol as the security parameter is increased. For formal definitions of both statistical and computational indistinguishability see Appendix A.
When talking about two indistinguishable families $\{\rho_{1}^{(k)}\}$ and $\{\rho_{2}^{(k)}\}$, if the parameter $k$ is implicit, we will just refer to them as $\rho_{1}$ and $\rho_{2}$ and use the following notation to denote indistinguishability:
$\rho_{1}\approx\rho_{2}$ $\quad\textnormal{for statistically indistinguishable};$
$\rho_{1}\approx^{(c)}\rho_{2}$ $\quad\textnormal{for computationally indistinguishable}.$
Additionally, in this work we consider protocols that can abort if certain conditions are satisfied. Mathematically, it is useful to consider the state of the aborted protocol as the zero operator. This means that events that trigger the protocol to abort are described as trace-decreasing operations, and hence, the operator representing the associated system at the end of the protocol is, in general, not normalized. The probability of the protocol finishing successfully is given then by the trace of the final state of the output registers. Note that the above definitions of indistinguishability can be naturally extended to non-normalized operators since the outcomes of a quantum program can always be represented by the outcomes of a POVM $\{F_{i}\}$, whose probabilities are given by Tr[$\Tr[F_{i}\rho]$, which is a well defined quantity even for non-normalized $\rho$.
Definition 2.1.
@@ 3 The protocol
In order to facilitate the finite-key security analysis, the description of $\pi_{\textnormal{QROT}}$ features two statistical tolerance parameters, denoted as $\delta_{1},\delta_{2}$. The role of $\delta_{1}$ is to account for the error in the estimation of the Qubit Error Rate (QBER), while the role of $\delta_{2}$ is to account for the small variations in the frequency of outcomes of $50/50$ events. These parameters can be ignored (set to zero) when considering very large values of $N_{0}$.
In the following description of the protocol we use the common conjugate coding notation used in BB84-based protocols. The bit values $0,1$ the denote the computational and Hadamard bases for qubit Hilbert spaces, respectively. For added clarity, we use the superscripts $A$ and $B$ to respectively denote Alice and Bob. Additionally, we use variable $x$ to denote measurement outcomes and $\theta$ to denote measurement bases (e.g. the pair $(\theta^{A}_{i},x^{A}_{i})$ denotes that Alice measured her $i$-th subsystem in the $\theta^{A}_{i}$ basis and obtained $x^{A}_{i}$ as the outcome). We use $|{\Phi^{+}}\rangle$} to denote the Bell state $\frac{1}{\sqrt{2}}({|{00}\rangle}+{|{11}\rangle})$. Finally, we will use the relative (or normalized) Hamming weight function $r_{H}:\{0,1\}^{n}\rightarrow[0,1]$ defined for any $x=(x_{1},\ldots,x_{n})$ as
@@
  • Parameter estimation sample ratio $0<\alpha<1$
@@
  • Maximum qubit error rate $0\leq p_{\text{max}}\leq 1/2$
@@
3.1 Security and performance of the main protocol
@@
A high-level proof of Theorem 3.1, including the derivation of the security bounds from Lemmas 3.1 and 3.2 can be found in Section 4 and further details can be found in Appendix B. The security of $\pi_{\textnormal{QROT}}$ is given by its main security parameter $N_{0}$, as well as the security parameters of the underlying commitment and IR schemes $k$ and $k^{\prime}$, respectively. These values can be computed for the statistical security features of the protocol and are given by the following lemmas:
Lemma 3.1.
@@
$\varepsilon^{\prime}$ $=\sqrt{2}\left(e^{-\frac{1}{2}(1-\alpha)^{2}N_{\textnormal{test}}\delta^{2}_{1}}+e^{-\frac{1}{2}N_{\textnormal{check}}\delta^{2}_{1}}\right)^{\frac{1}{2}}+e^{-D_{KL}(\frac{1}{2}-\delta_{2}|\frac{1}{2})(1-\alpha)N_{0}}+\varepsilon_{\textnormal{bind}}(k)$ (13)
@@
For the purposes of this analysis, we assume that the commitment and IR schemes, as well as their security parameters $k,k^{\prime}$, are appropriately chosen to satisfy the desired security level and we focus on the dependence of $\varepsilon_{\max}$ on the channel error rate, characterized by the parameter $p_{\max}$, and the number of quantum signals $N_{0}$. We are also interested in a quantity known as the secret key rate $R_{\textnormal{key}}$. For given values of $N_{0}$, $\alpha$, $\delta_{1}$, $\delta_{2}$, $p_{\max}$, and $\varepsilon_{\max}$, let $n_{\max}$ be the largest number for which the associated $n_{\max}$-bit ROT has at least security $\varepsilon_{\max}$, then
@@
Figure 2: Maximum key rate output $\frac{n}{N_{0}}$ versus error rate $p_{\max}$. The blue line represents the upper bound for the key rate, when $N_{0}\rightarrow\infty$, $\alpha,\delta_{1},\delta_{2}$ are taken to be 0 and $f=1$. The orange line represents a more typical case with $\alpha=0.35$, $\delta_{1}=0.01,\delta_{2}=0.025$, and $f=1.2$.
Figure 3: Maximum key rate behaviour as a function of $N_{0}$ for different security levels. Parameter values used are $\alpha=0.35;\delta_{1}=$9.20\text{\times}{10}^{-3}$;\delta_{2}=$3.00\text{\times}{10}^{-3}$;p_{\textnormal{max}}=0.01;f=1.2$.
@@
3.2 Experimental implementation performance
@@
where $N_{\textnormal{crit}}$ is computed using the optimal values of $\alpha,\delta_{1},\delta_{2}$ for the respective error rate $p_{\max}$ and undetected multi-photon rate $p_{\textnormal{multi}}$ associated to $P$, assuming perfectly efficient information reconciliation, $f=1$ (see Section 5 for the details on the implementation and its security). As seen in Figure 5, for this implementation, the additional coincidence rate gained by increasing $P$ is not enough to compensate for the induced increased error. This result is not immediately obvious, as $N_{\textnormal{crit}}$ does not depend explicitly on $p_{\max}$. The decrease in performance comes from the fact that increasing $p_{\max}$ limits the values that $\delta_{1}$ can have while maintaining positive key rates. This restriction on the values of $\delta_{1}$ ultimately results in an increase in $N_{\textnormal{crit}}$ and therefore, a reduction on $R_{OT}$.
@@
Table 1 shows an example of the performance of the protocol in a real-world implementation using the data from the experimental setup. For the commit/open phase, the weakly-interactive string commitment protocol introduced in [35] was implemented using the BLAKE3 hash function algorithm as a one-way function. For the post-processing phase, a low density parity check (LDPC) code was used for IR, and random binary matrices were used to implement the universal hash family for privacy amplification. We evaluated the performance by the number of $128$-bit ROT instances able to be completed per second (It is worth noting that, using a Mac mini M1 2020 16GB computer, the post-processing throughput was enough to handle all the data from the experiment, the bottleneck being the quantum signal generation rate).
@@
Security level $\varepsilon_{\textnormal{max}}$ $1.91\text{\times}{10}^{-8}$
Cost in quantum signals $N_{0}$ $5.86\text{\times}{10}^{6}$
@@
Statistical parameter 1 $\delta_{1}$ $9.00\text{\times}{10}^{-3}$
Statistical parameter 2 $\delta_{2}$ $3\text{\times}{10}^{-3}$
@@
Max allowed multi-photon rate $p_{\textnormal{multi}}$ $3.67\text{\times}{10}^{-3}$
@@
4 Security Analysis
In this section, we prove the main security result, which relates the overall security of the protocol as a function of its parameters $N_{0},\alpha,\delta_{1}$, and $\delta_{2}$ in Theorem 3.1. For clarity of presentation, we have compacted some of the properties into lemmas, for which detailed proofs can be found in Appendix B. Definitions and properties of entropic quantities can be found in Appendix A
4.1 Correctness
@@
Recall that we model the aborted state of the protocol as the zero operator. This way, whenever we have a mixture of states, some of which trigger aborting and some that do not, the abort operation removes the events that trigger it from the mixture, effectively reducing its trace by the probability of aborting. There are three instances where the protocol can abort: first during Step (7) if the estimated qubit error rate is larger than $p_{\max}$; the second one is during Step (9) if Bob does not have enough (mis)matching bases to construct the sets $I_{0},I_{1}$; and finally during Step (11) if the IR verification fails. The probability of aborting in Steps (7) and (11) depends on the particular transformation that the states undergo when being sent from Alice’s to Bob’s laboratory, about which we make no assumptions. We can group these three abort events and denote by $p_{\textnormal{abort}}$ the probability of the protocol aborting by the end of Step (11). The state at this point can be written as $(1-p_{\textnormal{abort}})\rho^{\top}$, where $\rho^{\top}$ represents the normalized state conditioned that the protocol has not aborted by this point. As Lemma 4.1 states, the verifiability property of the Information Reconciliation scheme guarantees that the states that “survive” past Step (11) have the property that Bob’s corrected string $y^{B}$ is the same as Alice’s outcome string $x^{A}_{I_{0}}$, which is uniformly distributed.
Lemma 4.1.
@@
During Step (12) universal hashing is used in both $x^{A}_{I_{0}}$ and $x^{A}_{I_{1}}$ to obtain $m_{c}$ and $m_{\bar{c}}$. Because Eq. (18) describes a state for which the $X^{A}_{I_{0}},X^{A}_{I_{1}}$, and $C$ subsystems are independent and uniformly distributed, it follows from Lemma 4.1 that
@@
Finally, using the quantum leftover hash Lemma A.4 twice (once for $m_{0}$ and $m_{1}$) with the corresponding entropy terms given by Eq. (4.1), together with Lemma A.1 (1), we conclude that the state $\rho^{(\textnormal{out})}_{M_{0},M_{1},C,M_{C}}$ of the output systems after the post processing phase satisfies (substituting $p_{\textnormal{succ}}=1-p_{\textnormal{abort}}$)
@@
4.2 Security against dishonest sender